sixty-two feet of fencing was used to enclose a rectangular pen that was 11 feet wide. How long was the pen?
Answer by checkley71(8403)

PERIMETERIS = 2 TIMES LENGTH + 2 TIMES WIDTH THUS
62=2*L+2*11
62=2L+22
62-22=2L
40=2L
40/2=L
20=LENGTH
PROOF
2*11+2*20=62
22+40=62
62=62